The sum of Mary's age and her mother's ages is 36. Mary's mother is five times older than Mary. How old is each person?
Answer by checkley77(12844)

x+5x=36
6x=36
x=36/6
x=6 age of Mary.
5*6=30 age of mom.
Proof:
6+30=36
36=36